OpenStudy (anonymous):

need help with following exression in factored form using the greastest common factor with 50a3 + 10a2

OpenStudy (anonymous):

well @Mehek14

Mehek (mehek14):

they have in common \(10a^2\)

Mehek (mehek14):

\(10a^2(?+?)\)

Mehek (mehek14):

divide \(\dfrac{50a^3}{10a^2}=5a\)

Mehek (mehek14):

divide \(\dfrac{10a^2}{10a^2}=1\)

Mehek (mehek14):

\(10a^2(5a+1)\)

OpenStudy (anonymous):

so its all about divide it

OpenStudy (anonymous):

thanks

Mehek (mehek14):

basically and np